OpenStudy (shiree19):

Can someone please help me solve this problem? It's about inequalities in algebra. I know how to do these it's just this one is a bit confusing for me.

OpenStudy (shiree19):

Here is the problem: \[q + 12 - 2(q - 22) > 0\]

OpenStudy (anonymous):

\[q >55\]

OpenStudy (shiree19):

Okay. But how did you get that?

OpenStudy (shiree19):

I would like to be able to do this kind of problem on my own in the future that's why I'm asking.

OpenStudy (anonymous):

it's wrong..

OpenStudy (shiree19):

Ok. Then what is the correct answer and how do you get it?

OpenStudy (anonymous):

@Shiree19 q + 12 - 2 (q - 22 ) > 0 simplify both sides -q + 56 > 0 subtract 56 from both sides -q + 56 - 56 > 0 - 56 - q > - 56 divide both sides by negative 1 -q/-1 > -56/-1 and you get q < 56 for the answer

OpenStudy (shiree19):

@cutelilcigarette Okay. (this will sound like I'm an idiot) But how did you simplify the left side?

OpenStudy (anonymous):

\[q+12-2(q-22)>0 first you multiple -2 by q then you multiply -2 by -22 then subtract 44 from 0 then subtract 12 from 0 then you then you take -2q and add 1q so that makes a -1q so then you -1q and devide by -56\]

OpenStudy (shiree19):

ok. So I do \[-2 \times q = -2q\] then I do \[-2q \times 22 = -44q\] than I do \[-44 - 12 = -56\] So that's -q + 56?

OpenStudy (shiree19):

Is that right? How do you get a negative q?

OpenStudy (anonymous):

-2q +1q

OpenStudy (shiree19):

oh ok. Thanks so much!